OpenStudy (anonymous):

1. Simplify 5^-1(3^-2)

OpenStudy (eujc21):

\[5^{-1^{3^{-2}}} \] or \[5^{-1}*3^{-2}\]

OpenStudy (anonymous):

5^-1 * 3^-2

OpenStudy (eujc21):

so.... what is \[5^{-1} == \frac{1}{5}\]

OpenStudy (anonymous):

Thats what I got for that. For my answer i got 15^2. Right?

OpenStudy (eujc21):

\[3^{-2} == \frac{1}{3^{2}} == \frac{1}{9}\] and finally \[\frac{1}{5}*\frac{1}{9}=? \]

OpenStudy (anonymous):

Ohh I see how you did that. So 1/45? Im bad at math.

OpenStudy (eujc21):

Yeap, it's best to do math step by step, eventually the shortcuts are engraved in your brain.

OpenStudy (anonymous):

Yeah I agree. Thanks. I now know what you did there.